OpenStudy (anonymous):

Given ABCD, AC = 28, and AE = 3x - 7, find the value of x.

myininaya (myininaya):

I don't know what ABCD is?

OpenStudy (anonymous):

this is ABCD

myininaya (myininaya):

assuming it is a parallelogram, those diagonals bisect each other. So then AE would be half of AC. like AE+AE=AC 2AE=AC...

OpenStudy (anonymous):

So does 14 sound right?

myininaya (myininaya):

idk did you do 2AE first?

myininaya (myininaya):

what is 2(3x-7)?

OpenStudy (anonymous):

-1

myininaya (myininaya):

how did you get -1 from 2(3x-7)?

myininaya (myininaya):

x is a variable is varies so 2(3x-7) will also vary

myininaya (myininaya):

2(3x)=? 2(-7)=?

myininaya (myininaya):

\[2(3x+7)=2(3x+-7)=2(3x)+2(-7)=\]

myininaya (myininaya):

what is 2(3)?